OpenStudy (anonymous):

A chemist needs 0.0700 mol of potassium for a reaction. What mass of potassium should the chemist use? 2.737 g K 2.74 g K 39.1 g K

OpenStudy (rushwr):

moles = mass divided by molar mass \[moles (n) = \frac{ mass (m) }{Molar mass (M) }\] So here we know the no. of moles and the molar mass of Potassium which is 39.0983 g per moles if moles = mass divided by molar mass then , mass = moles * molar mass therefore mass = 0.07 * 39.0983= 2.737gK

OpenStudy (anonymous):

ok that makes more sense thank you